(sin 70 +cos 40)/(cos 70+sin 40)=
Answer by math-vortex(648)

Hi, there--

THE PROBLEM:
Evaluate (sin 70+cos 40)/(cos 70+sin 40)

A SOLUTION:

(sin 70+cos 40)/(cos 70+sin 40) = (0.9397+0.7660)/(0.3420+0.6428) = 1.7320

This is an approximate answer, rounded to 4 decimal places.


Hope this helps! Feel free to email if you have any questions about the solution.

Good luck with your math,

Mrs. F
math.in.the.vortex@gmail.com